'A(aq) 2 B(aq)-----> AB 2 (aq) suppose that 40ml of .0116 M A(aq) is mixed with 50ml of .0202 M B(aq) and after the reaction .000375 moles of AB2 were obtained calculate the moles of A and B before the reaction
1
Expert's answer
2012-03-13T07:54:14-0400
Moles of A:

n(A) = 40*10-3 dm3 * 0.0116 M = 0.000464 mol

Moles of
B:

n(B) = 50*10-3 dm3 * 0.0202 M = 0.00101 mol